Find Z[n(n – 1) (n – 2)].


1
Expert's answer
2021-07-12T13:58:11-0400

The Z-transform of following function will be


"Z[n(n-1)(n-2)]=Z[n(n^{2}-3n+2)]=Z[n^{3}-3n^{2}+2n]"



"=Z[n^{3}]-3Z[n^{2}]+2Z[n]"


"=" We know that "Z" transform of "n^{3}" is given by


"Z[n^{3}]=\dfrac{z^{3}+4z^{2}+z}{(z-1)^{4}}"


"Z[n^{2}]=\dfrac{z^{2}+z}{(z-1)^{3}}"


"Z[n]=\dfrac{z}{(z-1)^{2}}"



Now putting these values in above equation , we get


"=" "\dfrac{z^{3}+4z^{2}+z}{(z-1)^{4}}+\dfrac{z^{2}+z}{(z-1)^{3}}+\dfrac{z}{(z-1)^{2}}"